Following is information on the price per share and the dividend for a sample of 30 companies.

Obs Price Dividend
1 20.21 3.14
2 22.22 3.36
3 31.60 0.46
4 33.78 7.99
5 36.07 0.77
6 36.33 8.46
7 36.37 7.62
8 38.20 8.03
9 39.06 6.33
10 39.86 7.96
11 43.65 8.95
12 49.29 9.61
13 53.94 11.11
14 54.62 13.28
15 55.31 10.22
16 57.27 9.53
17 57.61 12.60
18 58.51 10.43
19 59.72 7.97
20 60.81 9.19
21 64.22 16.50
22 64.87 16.10
23 64.95 13.76
24 65.16 10.54
25 66.64 21.15
26 68.39 14.30
27 69.77 24.42
28 75.11 11.54
29 78.12 17.65
30 80.21 17.36

a. Calculate the regression equation that predicts price per share based on the annual dividend (round answer to 4 decimal places)

Price = ____ + _____ dividend

b1. State the null and alternate hypotheses

a. H0: β = 0 and H1: β ≠ 0

b. H0: β = 0 and H1: β = 0

c. H0: β ≠ 0 and H1: β ≠ 0

d. H0: β ≤ 0 and H1: β > 0

b2. State the decision rule. Use the 0.05 significance level (round answer to 3 decimal places)

Reject H0 if t is not between ± ______

b3. Compute the value of the test statistic (round answer to 4 decimal places)

The test statistic is ______

b4. What is your decision regarding H0?

a. do not reject

b. reject

b5. Can we conclude the slope is not equal to zero?

a. yes

b. no

c. Determine the coefficient of determination? (round answer 4 decimal places)

R2 ______

d1. Determine the correlation coefficient (round answer 4 decimal places)

Correlation coefficient ______

d2. Can you conclude that it is greater than 0 using the 0.05 significance level?

______ H0. We _____ (a. can, b. cannot) conclude that the correlation in the population is _____ (a. greater than, b. less than) Zero.

e. If the dividend is $10, what is the predicted price per share? (round 4 decimal places)

Price ______

f. What is the 95% prediction interval of price per share if the dividend is $10? (round 4 decimal places)

[ _____ , ____ ]
Answer #1

Answer:

Here we use Excel to fit a regression line. We use regression option in data analysis from Excel.

Here the response variable y is Price the explanatory variable x is Dividend.

The output of the regression can be given as

D E F G H L M N SUMMARY OUTPUT Regression Statistics Multiple R 0.811358445 R Square 0.658302526 Adjusted R Square 0.64609904

Using the above regression output, segression equation can be given The Price = 27.0154 + 2.4082 Dividend. line Now the test

Decision : Reject Reject Ho if I teall > #0-2,412) here x = 0.05 = 0.025 2 do tem-2,212) + 30-2, 0.025) = 728,0·025) = 2.0484